Find the energy transferred in 1 h by conduction through a concrete wall 2 m high, 3.65 m long, and 0.20 m thick if one side of the wall is held at 20∘C and the other side is at 5∘C.

1
Expert's answer
2022-03-13T18:48:57-0400

The Fourier law of heat conductivity says

"Q=P\tau=k\frac{\Delta T}{\Delta h}A\tau"

"Q=1.9*\frac{20-5}{0.20}*2*3.65*1*3600\\
=3.7*10^6\:\rm J"